OpenStudy (anonymous):

Solve AX=0 for X = [x] [y] [z]

OpenStudy (anonymous):

\[X = \left(\begin{matrix}x \\ y \\ z\end{matrix}\right)\]

OpenStudy (zzr0ck3r):

augment the matrix with 0 vector and row reduce

OpenStudy (anonymous):

A = \[A =\left[\begin{matrix}1 & 1 &1 \\ 1 & 2&1\\2&3&2\end{matrix}\right]\]

OpenStudy (anonymous):

augment? i'm not familiar with that term

OpenStudy (anonymous):

so do i dot product X with A?

OpenStudy (zzr0ck3r):

can you put the matrix into row reduced echelon form?

OpenStudy (anonymous):

sorry, i'm not familiar with that either

OpenStudy (anonymous):

well 0, but where did 3x = 0 come from?

OpenStudy (anonymous):

do i dot product A with X, then set the corresponding rows equal to 0?

OpenStudy (zzr0ck3r):

forget I said that...what are you doing in class?

OpenStudy (anonymous):

we just learned eigenvalues and vectors, but i believe this is right before that section

OpenStudy (zzr0ck3r):

x+y+z=0 x+2y+z=0 2x+3y+2z=0

OpenStudy (anonymous):

okay, so it's basically a dot product, then i have to solve the system of equations

OpenStudy (zzr0ck3r):

you have a system with three unknowns and three equations you can solve this as you would any system. but there are nice tricks for solving these, but would be hard to explain on here. google "row reduced echelon form".

OpenStudy (zzr0ck3r):

so I take it you are in a dif eq class and not linear algebra?

OpenStudy (anonymous):

yes, this is diff eq

OpenStudy (zzr0ck3r):

gotcha. yeah just solve like any system...

OpenStudy (anonymous):

okay, thank you

OpenStudy (zzr0ck3r):

np